OpenStudy (ascendant):

Calculate the number of days during the month of October that were snowy, rainy, sunny, and cloudy. (Hint: There are 31 days in October) Cloudy-19% Sunny-13 % Rainy- 23% Snowy- 45%

OpenStudy (ascendant):

I know what the percentage is as a decimal, I just don't know how to covert a decimal into a whole number.

OpenStudy (mathmale):

Convert each given percentage to a ratio and then multiply 31 days by it. Try this now, please.

OpenStudy (ascendant):

Sorry, but how do you convert a percentage into a ratio?

OpenStudy (skullpatrol):

$$\huge \dfrac{x}{100}=\dfrac{a}{b}$$ $$\huge \dfrac{x}{100}=\dfrac{a}{31}$$

OpenStudy (skullpatrol):

$$\huge \dfrac{x}{100}=x\text{%}$$

OpenStudy (skullpatrol):

$$\huge \dfrac{45}{100}=\dfrac{a}{31}$$

OpenStudy (skullpatrol):

Can you solve for a?

OpenStudy (ascendant):

Yes, Thank you for your help!

OpenStudy (skullpatrol):

Thanks for trying to understand :-)

OpenStudy (ascendant):

You're Welcome

OpenStudy (skullpatrol):

Just remember percent means per 100 or 1/100

OpenStudy (skullpatrol):

$$\huge x\text{%}=\dfrac{x}{100}=\dfrac{a}{b}$$

OpenStudy (ascendant):

I'm sorry but how do you solve for A?

OpenStudy (skullpatrol):

Multiply both sides by 31.

OpenStudy (skullpatrol):

$$\huge \dfrac{x}{100}=\dfrac{a}{b}$$ $$\huge \dfrac{45}{100}=\dfrac{a}{31}$$ $$\huge31• \dfrac{45}{100}=\dfrac{a}{31} •31$$